OpenStudy (anonymous):

Simplify the expression and eliminate any negative exponent(s). Assume that y and z denote positive numbers. (16y^-2z^4)^1/2 over (8y^3z^-6)^-1/3

OpenStudy (anonymous):

\[\frac{(16y^{-2}z^4)^\frac{1}{2}}{(8y^3z^{-6})^{\frac{1}{3}}}\]

OpenStudy (anonymous):

ok lets get rid of the fractional exponents first by multiplying every exponent in the numerator by 1/2 and every one in the denominator by 1/3 also know that \[16^{\frac{1}{2}}=\sqrt{16}=4\]\]

OpenStudy (anonymous):

and likewise \[8^{\frac{1}{3}}=\sqrt[3]{8}=2\]

OpenStudy (anonymous):

so the numerator will be \[4y^{-1}z^2\] and the denominator will be \[2yz^{-3}\]

OpenStudy (anonymous):

our fraction is now \[\frac{4y^{-1}z^2}{2yz^{-2}}\]

OpenStudy (anonymous):

ok whats next

OpenStudy (anonymous):

to make the negative exponents positive, if it is up bring it down and if it is down bring it up so you get \[\frac{4z^2z^2}{2yy}\]which gives \[\frac{2z^4}{y^2}\]

OpenStudy (anonymous):

i also of course divided 4 by 2 to get 2 in the numerator

OpenStudy (anonymous):

gotcha. whats the final

OpenStudy (anonymous):

notice that the \[z^{-2}\] in the denominator came up to the numerator as \[z^2\] and the \[y^{-1}\] in the numerator came to the denominator as \[y\]

OpenStudy (anonymous):

"final answer" is \[\frac{2z^4}{y^2}\]

OpenStudy (anonymous):

\[\color{green}{\text{hello karnak}}\]